# Gamma-ray burst Jet Progenitor: long vs short bursts by angular momentum and mass ratio

### v1

@@ 7.3 Long vs short GRBs
    The jet and sphere size ratio $R_{\gamma}^{j}/R_{\gamma}\sim{\mathcal{O}}(10^{-1})$. It implies that long bursts duration $T_{90}\sim R_{\gamma}/(2c\Gamma^{2})$ are about 10 times lager than short bursts $T^{j}_{90}\sim R^{j}_{\gamma}/(2c\Gamma_{j}^{2})$. We assume hydrodynamic expansion Lorentz factors $\Gamma/\Gamma_{j}\sim{\mathcal{O}}(1)$, considering the obtained result $E^{j}_{\gamma}/E_{\gamma}\sim{\mathcal{O}}(1)$ in Table 1. It is qualitatively consistent with two classes of long (soft) and short (hard) bursts’ data scattering on the $E_{p}$-$T_{90}$ plane, see Figures 1 and 6 (up-right) of the reference [25]. However, the separatrix (7.1) of this simplified model cannot infer the separatrix $T_{90}\approx 1\sim 2$ seconds between long GRBs (massive core collapses) and short GRBs (binary coalescence).
@@
    The trends of numerical results in Table 1 imply possibly positive or negative correlations between the jet angle $\theta_{\rm jet}$ and burst duration $T_{90}^{j}$, spectral peak $E_{p}^{j}$, total energy $E_{\gamma}^{j}$, luminosity $L_{\gamma}^{j}$. Short GRBs should be more jetting than long GRBs.
These insights underscore the importance of considering rotational effects and jet dynamics in studying GRB spectra and their correlations.
8 Conclusion and remarks
@@
To understand GRBs’ jet progenitors, we adopt an adiabatic approximation and a simplified model to study photon-pair jets generated by the gravitational collapse of slowly rotating cores and binary systems. Our results qualitatively show photon-pair jets features: (i) the jet angle $\theta_{\rm jet}$ relates to progenitor angular momenta and masses ratio $J/M$; (ii) variation trends of GRBs’ peak spectra, total energy and luminosity in terms of jet angle, positive and negative correlations among them; (iii) the distinction between short and long GRBs’ properties is due to the different angular momentum of systems that undergo gravitational processes of binary coalescence or massive core collapse.
More theoretical studies and numerical simulations of massive rotating core collapse and binary coalescence are necessarily required to better understand the dynamics of massive rotating core collapse and binary coalescence. These studies will help refine our understanding of the photon-pair jet formation process and the resultant GRB properties.
@@
The data analyses on possible correlations of GRBs’ jet energetic, geometric and time-duration features are inviting. Systematic analyses of GRB data are required to explore possible correlations between jet energetics, geometric features, and time-duration characteristics. Understanding these correlations will provide deeper insights into the jet progenitor systems and the physical mechanisms driving GRBs.
In summary, our simplified model provides a qualitative framework to understand the generation of photon-pair jets from gravitationally collapsing systems. It highlights the significant role of progenitor angular momentum in determining the jet characteristics and, consequently, the observable properties of GRBs. Further theoretical and observational studies are crucial to validate and expand upon these findings, paving the way for a more comprehensive understanding of GRB jet progenitors and their evolution.

### v2

@@ 9.2 Long vs short GRBs and jet angles
    The collimatied and sphere fireball sizes’ ratio $R_{\gamma}^{j}/R_{\gamma}\sim{\mathcal{O}}(10^{-1})$. It implies that long bursts duration $T_{90}\sim R_{\gamma}/(2c\Gamma^{2}_{0})$ (8.5) are about 10 times lager than short bursts $T^{j}_{90}\sim R^{j}_{\gamma}/(2c\Gamma_{0j}^{2})$. Here, we assume hydrodynamic expansion ejecta Lorentz factors $\Gamma_{0}/\Gamma_{0j}\sim{\mathcal{O}}(1)$, considering the obtained result $E^{j}_{\gamma}/E_{\gamma}\sim{\mathcal{O}}(1)$ in Table 1 and assuming ejecta mass $M_{\rm ej}$ is the same order of magnitude in spherical and collimated cases, i.e., $M^{j}_{\rm ej}/M_{\rm ej}\sim{\mathcal{O}}(1)$. The ratio $T^{j}_{90}/T_{90}\sim{\mathcal{O}}(10^{-1})$ is qualitatively consistent with two distinct classes of long (soft) and short (hard) bursts’ time-duration data scattering on the $E_{p}$-$T_{90}$ plane, see Figures 1 and 6 (up-right) of the reference [47].
@@
    The trends of numerical results as increasing $J/M$ in Table 1 imply possibly positive or negative correlations between the collimated fireball angle $\theta_{\rm coll}$, burst duration $T_{90}^{j}$, spectral peak $E_{p}^{j}$, total energy $E_{\gamma}^{j}$, luminosity $L_{\gamma}^{j}$. These trends indicate that short GRBs (hard photon spectrum) tend to be emitted from a narrow collimated fireball compared to long GRBs (soft photon spectrum) preferentially from a nearly spherical fireball. In other words, short GRBs from binary coalescences of large angular momenta $J/M$ are more jetting (smaller $\theta_{\rm coll}\sim\theta_{\rm jet}$) than long GRBs from rotating core collapses of small angular momenta. Short GRBs originate from more collimated fireballs with large angular momentum $J_{\gamma}/M$. Long GRBs originate from less collimated fireballs with small angular momentum $J_{\gamma}/M$.
  6. (vi)
    The above observations on the trends vs angular momentum $J/M$ and Sec. 9.3 discussions on the separatrix $(J/M)_{\rm sepa}$ indicate that short GRB progenitors possess larger $J/M$ values than long GRB progenitors. The separatrix (9.1) of this oversimplified model cannot infer the separatrix $T_{90}\approx 1\sim 2$ seconds between long GRBs (massive core collapses) and short GRBs (binary coalescence). However, following the discussions on the viewpoint (iv) above, we find that $R^{j}_{\gamma}/R_{\gamma}=\tau^{j}_{\gamma}/\tau_{\gamma}\approx 10^{-1}$ corresponds to the line $a^{2}\Omega_{c}^{2}\approx 0.02$ indicated by $\Rightarrow$” in Table 1. It infers collimated fireball angles $\theta_{\rm coll}<5^{\circ}$, thus observed jet angles $\theta^{*}_{\rm jet}<5^{\circ}$ for the cases of short GRBs originating from massive binary coalescence. The separatrix of GRB jet angles should be in short and long GRB data.
We have not found inconsistencies between the theoretical scenario and observational phenomena, although the adopted model and approximation are oversimplified for complex GRB progenitor systems. However, the model adopted is still far from a realistic application in analyzing GRB progenitor properties and their consequences. Despite these weaknesses, these scenario insights underscore the importance of considering the angular momentum of rotating systems in collapses and collimated fireball dynamics in studying GRBs’ spectrum, energy, luminosity, duration, jetting angle, and their correlations.
9.3 Angular momentum separatrix of core collapses and binary mergers
In addition, the distinct separation between long and short GRB classes is evident in observations. We attempt to give a possible explanation of this phenomenon consistently with our arguments on the properties of GRBs jet progenitors varying with its $J/M$ in Sec. 9.1. For larger $J/M$ values and rotational repelling energies $U_{c}/|U|$ (5.33), a massive rotating core cannot maintain the equilibrium conditions (5.15) with given material binding energy $\lambda$. Thus, rotating stellar cores split. In this case, the most probable configuration is a binary system. This condition indicates two distinct classes: massive rotating core collapses with small $J/M$ values and binary mergers with large $J/M$ values.
We speculate that the separatrix between two distinct classes occurs when the equilibrium condition (5.10) or (5.15) breaks down. Namely, the centrifugal potential $\phi_{c}$ and heat function $h$ are too large to be balanced by the gravitational binding potential $\phi$ and the material binding energy $\lambda$. Note that the $\lambda$ value should not depend on whether the system is a single core or a binary object at the nuclear density.
Considering the particular case that the heat function $h$ balances the gravitational binding potential $\phi$ in the equilibrium condition (5.10), we find the critical centrifugal potential balanced by material binding potential energy $\lambda$,
$(\phi_{c})_{\rm sepa}\approx\frac{(v_{s}^{c})^{2}}{\gamma-1}\frac{\lambda}{6}r^{2}_{s}\quad\Rightarrow\quad(a^{2}\Omega^{2})_{\rm sepa}\sim\frac{\lambda}{3(\gamma-1)},$ (9.1)
where $\phi_{c}\approx\frac{(v_{s}^{c})^{2}}{2}r^{2}_{s}a^{2}\Omega_{c}^{2}$ (5.26) by setting $r=r_{s}$ and $\sin^{2}\theta=1$. The critical condition (9.1) indicates a separatrix $(J/M)_{\rm sepa}$ via the angular-momentum relation (5.30) or a separatrix $(U_{c}/|U|)_{\rm sepa}$ via the rotation-energy relation (5.33). This separatrix implies the following two separate situations:
  1. (a)
    When the material binding energy is larger than the repulsive rotation energy, all materials bind together and undergo massive core collapses. This case is for slowly rotating cores $J/M\ll 1$ or $U_{c}/|U|\ll 1$. Such progenitors produce less collimated fireballs with small angular momentum $J_{\gamma}/M\ll 1$.
  2. (b)
    On the contrary, all materials cannot bind together as a single massive core in collapses when the systems have large $J/M$ or $U_{c}/|U|$ values, and the material binding energy is smaller than the repulsive rotation energy. Therefore, some gravitational binding energy is required to balance rotational energy. Such systems are very probably realised as binary systems undergoing gravitational mergers. Such progenitors produce more collimated fireballs with large angular momenta $J_{\gamma}/M$.
Although we are not able to calculate the separatrix, these discussions give an insight into the distinction between long and short GRBs’ classes relating to the angular momentum and mass ratios $J/M$ of massive rotating core collapses and binary coalescence. The realistic situation of separating two distinct GRB classes should be much more complex. Our theoretical model and analysis methods are too simplified and approximate for fast-rotating core collapses and binary coalescence. Therefore, our physical explanations of jet angle correlation, short GRBs and long GRBs distinction in terms of $J/M$ are speculative. Further theoretical and empirical studies on these explanations are necessary and inviting.
10 Summary and remarks on GRB light angular momenta
@@
In summary, to understand the physics of progenitors generating highly collimated GRB events, we adopt a simplified model in the Newtonian approximation to qualitatively study photon-pair collimated fireballs generated by the gravitational collapse of slowly rotating massive cores and binary systems. Our qualitative studies infer the following photon-pair collimated fireball features: (i) collimated fireballs acquire nontrivial angular momenta $J_{\gamma}$ from massive rotating systems (progenitors) in gravitational collapses and make ultra-relativistic jets due to the angular momentum $J_{\gamma}$ conservation along propagating direction; (ii) the progenitor angular momentum and mass ratio $J/M$ determines the collimated fireball angle $\theta_{\rm coll}$ relating to observed GRB jet angles; (iii) collimated fireball quantities ($\theta_{\rm coll}$, peak spectra, total energy, luminosity, and time scale) have positive and negative correlations among them, whose imprints on observational quantities can be examined by using a large sample of observed GRB data; (iv) the distinctions between short and long GRBs’ properties are due to the different ratios of the angular momenta and masses of systems that undergo gravitational processes of massive binary coalescence or core collapse. These highlight the significant role of progenitor angular momentum and mass ratios in determining the collimated fireball characteristics and, consequently, the observable properties of GRBs, paving the way for a more comprehensive understanding of GRB jet progenitors.
On the one hand, more theoretical studies and numerical simulations of massive rotating core collapse and binary coalescence are necessary to understand better how GRB progenitors generate collimated fireballs with nontrivial angular momenta along the rotating direction of collapse systems, and such collimated fireball outflows launch ultra-relativistic jets, propagating with angular momentum conservation along jet directions. These studies from the first principle will help refine our understanding of the dynamics of collimated fireball formation, ultra-relativistic jet launching and propagation, and resultant GRB properties.
@@
On the other hand, it is important to perform the data analyses on empirical correlations of short and long GRB data, connecting to collimated fireball characteristics: energetic, geometric, and temporal features. These data analyses will reveal the intrinsic correlations among collimated fireball characteristics, thus providing deeper insights into the jet progenitor systems and the physical mechanisms driving GRBs.
To end this article, we mention that emitted from an ultra-relativistic jet with nontrivial angular momentum $J_{\gamma}$ along the axis of propagation and propagating in external magnetic fields, GRB light should possibly carry nontrivial spin angular momentum (SAM) and orbital angular momentum (OAM). The former leads to circular wave polarization, and the latter makes the helical twisting wavefront. These issues could be interesting for GRB theoretical studies and observational analyses.